OpenStudy (konradzuse):

Bags of a certain brand of tortilla chips claim to have a net weight of 14 ounces. Net weights actually vary slightly from bag to bag and are Normally distributed with mean μ. A representative of a consumer advocate group wishes to see if there is any evidence that the mean net weight is less than advertised and so intends to test the hypotheses H0: μ = 14, Ha: μ < 14. To do this, he selects 16 bags of this brand at random and determines the net weight of each. He finds the sample mean to be = 13.88 and the sample standard deviation to be s = 0.24. Reference: Ref 16-4 Based on these data, A. we would reject H0 at significance level 0.10 but not at 0.05. B. we would reject H0 at significance level 0.05 but not at 0.025. C. we would reject H0 at significance level 0.025 but not at 0.01. D. we would reject H0 at significance level 0.01.

OpenStudy (konradzuse):

@jim_thompson5910

OpenStudy (konradzuse):

The only significance levels I know of are 0.010 and 0.05.. A?

OpenStudy (konradzuse):

or how do you solve for the levels?!

jimthompson5910 (jim_thompson5910):

Test Statistic: t = (xbar-mu)/(s/sqrt(n)) t = (13.88-14)/(0.24/sqrt(16)) t = -2

OpenStudy (konradzuse):

n = 16 or 14?

jimthompson5910 (jim_thompson5910):

there are 16 items in the sample, so n = 16

OpenStudy (konradzuse):

ic ic ok..

jimthompson5910 (jim_thompson5910):

P-Value: p = tcdf(-10,-2,15) = 0.03197 So you reject H0 if alpha = 0.05 or alpha = 0.10 http://www.wolframalpha.com/input/?i=tcdf%28-10%2C-2%2C15%29

OpenStudy (konradzuse):

SO... A!

jimthompson5910 (jim_thompson5910):

no

OpenStudy (konradzuse):

Then I'm confused.. :(

jimthompson5910 (jim_thompson5910):

you reject H0 if alpha = 0.05 or alpha = 0.10

OpenStudy (konradzuse):

Ic so A is false then. I'm not sure what Wolfram is exactly showing us....?

jimthompson5910 (jim_thompson5910):

it's showing the p value

OpenStudy (konradzuse):

Ic hmm...

OpenStudy (konradzuse):

So how do we test the other levels?

OpenStudy (konradzuse):

Damn I'm in trouble... 4 Q's in 30 mins..... :(

jimthompson5910 (jim_thompson5910):

"test the other levels"?

OpenStudy (konradzuse):

The Sig levels, for B, C and D?

jimthompson5910 (jim_thompson5910):

oh, the p-value is 0.03197 so it's only significant if alpha is 0.05 or 0.10 so you're rejecting H0 if alpha is 0.05 or 0.10

OpenStudy (konradzuse):

Oh, so B?

jimthompson5910 (jim_thompson5910):

but not at 0.025 since 0.03197 is not less than 0.025, so the answer is B

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (konradzuse):

:) NEXT Q GOGOG FAST!